# Empirical Bernstein and betting confidence intervals for randomized quasi-Monte Carlo

### v1

@@ 3 Asymptotic confidence interval comparisons
Our confidence intervals will use $R$ IID replicates that each use $n$ RQMC points in $[0,1]^{d}$. That takes $N=nR$ evaluations of $f$. Here we study how to get the narrowest confidence intervals with that budget of $N$ function evaluations.
@@
We know from WauRam24a that the widths of their predictable plug-in EBCI approaches that of some Bennett intervals described below, as $R\to\infty$ for fixed $n$. The same is true of their HBCI. Here we study the value of $n$ that optimizes the Bennett intervals.
As mentioned above, the tilde in $\tilde{O}$ hides powers of $\log(n)$. Those powers are present in error bounds for adversarially chosen integrands but do not seem to come up in real problems schl:2002 ; wherearethelogs . We will consider working models of the form
@@
when using $R$ IID observations that each have variance $\sigma^{2}_{n}$. An oracle that knew $\sigma^{2}_{n}$ could use this formula to select the best $n$ and $R$ for a confidence interval subject to a constraint $nR=N$. The predictable plug-in EBCI and the HBCI from WauRam24a do not assume a known variance for the $Y_{i}$. However those intervals attain the same asymptotic limiting width as $R\to\infty$ that an oracle would get from Bennett’s formula.
@@
If $\theta=1$, then $H(n)$ is minimized over $[1,\infty)$ at $n=1$.
@@
For scrambled Sobol’ points it is generally best to take $n$ to be a power of 2 Owe22a , which is not reflected in Theorem 1. We incorporate that restriction into some numerical examples in Section 4.
As noted above, when $\theta=1$, the best $n$ is $n=1$. In that case RQMC is the same as MC, using one uniformly distributed point in $[0,1]$. This result holds for any sampler that gets the ordinary MC rate, even if it has a favorable constant. In particular, any MC variance reduction techniques that do not change the convergence rate will not lead to using $n>1$. They will improve the asymptotic confidence interval width by reducing $\sigma_{0}$.
@@ 4 Finite sample evaluations
Here we present some finite sample evaluations of the EBCI (specifically the predictable plug-in one) and the HBCI, both using RQMC, for 95% confidence. We begin with finite $N$ and for some integrands with known RQMC variance, we find the $n$ that minimizes the interval widths from Bennett’s inequality. They are then the exact best $n$ for Bennett’s inequality which the HBCIs and EBCIs of WauRam24a match asymptotically. After that we present finite sample results on the EBCIs and HBCIs for a variety of integrands differing in smoothness and dimension. The HBCIs are usually narrower. Finally, we compare the empirically optimal $n$ for the HBCIs to the values that optimize Bennett’s inequality. The HCBI widths approach those from Bennett’s inequality for large $R$ and we find that the optimal $n$ for HBCI tend to be the same or smaller than the ones we get from Bennett’s inequality.
For our HBCI computations, we used $\theta=1/2$. The software we used comes from hedged_algo which has code to reproduce the figures in WauRam24a . It calls the HBCI and predictable plug-in EBCI algorithms from bet_algo . The scrambled Sobol’ points we used were from QMCPy QMCPy2020a .
@@ 4.1 Semi-empirical evaluations
The digital nets of Sobol’ Sob67 under the nested uniform scramble of Owe95 have a stratification property. The $n$ values of $x_{ij}$ for $i=1,\dots,n$ and any component $j=1,\dots,d$ have one value uniformly distributed in each interval $[(\ell-1)/n,\ell/n)$ for $\ell=1,\dots,n$ and those $n$ values are mutually independent. This allows us to work out the finite sample RQMC variances for some illustrative examples. The random linear scramble of Mat98 combined with a digital shift attains the same variance as the nested uniform scramble, but it has some dependencies among the stratified sample values. In both cases, the $d$ vectors $(x_{1j},\dots,x_{nj})\in\mathbb{R}^{n}$ are mutually independent.
We only consider values of $n$ that are powers of $2$ which is a best practice when using Sobol’ sequences Owe22a . The smooth error rate $\tilde{O}(n^{-3/2})$ for the RMSE is attainable along powers of $2$ but not for general $n$. An elegant argument due to Sobol’ sobo:1998 explains why. He notes that
@@
For $d=1$ and $f$ of bounded variation in the usual sense, we get $\operatorname{var}(\hat{\mu})\leqslant V_{\mathrm{HK}}(f)^{2}/n^{2}$. This holds because $D_{n}^{*}$ for the stratified sample cannot be larger than $1/n$. For illustration we choose a function with $V_{\mathrm{HK}}(f)=1$ for which we know exactly what the RQMC variance is for $n$ a power of $2$. That integrand is $f(x)=1\{x<1/3\}$. It is constant within $n-1$ of the intervals $[(\ell-1)/n,\ell/n)$. All of this function’s variation happens within just one of those intervals. By inspecting the base $2$ expansion $1/3=0.010101\cdots$ we can see that the remaining value of $f(x_{\ell})$ is $1$ or 00 and it equals $1$ with a probability that is either $1/3$ if $\log_{2}(n)$ is an odd integer or is $2/3$ if $\log_{2}(n)$ is an even integer. It follows that $\operatorname{var}(\hat{\mu}_{n})=2/(9n^{2})$.
Table 1 shows how the width minimizing value of $n$ increases with $N$ in the BVHK case with $f(x)=1\{x<1/3\}$. These are widths $W=2H^{\mathrm{Ben}}(n,N/n)$ using equation (11), not half widths. Some are larger than $1$, because the Bennett width does not reflect that we would intersect the intervals with $[0,1]$. The increase in $n$ is quite slow. By $N=1024$ the best RQMC sample size is only $n=8$ and this remains the optimal sample size up to $N=4096$. The value $N^{2/3}W$ very quickly becomes constant. For non-integer $\theta$ (not shown here) the scaled widths fluctuation within a finite range instead of approaching a constant.
@@
$\log_{2}(N)$ $N$ $n$ $W$ $N^{2/3}W$
00 1 1 5.02e$+$00 5.020114
04 16 2 7.60e--01 4.826379
07 128 4 1.90e--01 4.826379
10 1,024 8 4.75e--02 4.826379
13 8,192 16 1.19e--02 4.826379
16 65,536 32 2.97e--03 4.826379
19 524,288 64 7.42e--04 4.826379
22 4,194,304 128 1.86e--04 4.826379
25 33,554,432 256 4.64e--05 4.826379
28 268,435,456 512 1.16e--05 4.826379
Table 1: As $N$ varies for $f(x)=1\{x<1/3\}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{2/3}W$. Each row corresponds to a value of $N$ for which $n$ has increased.
Our second example is $f(x)=xe^{x-1}$. This function is bounded between 00 and $1$. We chose it because it is smooth enough to have $\theta=3$ while also not a polynomial or an antisymmetric function which would make the problem artificially easy for some methods. Because of the stratification, we know that
@@
Table 2 shows how the optimal sample sizes $n$ grows with $N$ for our smooth example $f(x)=xe^{x-1}$. An RQMC sample size of $n=8$ becomes optimal for $N=4{,}096$ and remains optimal up to $N=32{,}768$. The values $N^{3/4}W$ settle down rapidly as $N$ increases. Consistent with the results of Section 3, the values $n$ grow at a slower rate with $N$ than in the BVHK case while the widths $W$ decrease at a faster rate.
$\log_{2}(N)$ $N$ $n$ $W$ $N^{3/4}W$
0 1 1 4.00e$+$00 4.003728
4 16 2 5.17e--01 4.138086
8 256 4 6.52e--02 4.175717
12 4,096 8 8.17e--03 4.185407
16 65,536 16 1.02e--03 4.187848
20 1,048,576 32 1.28e--04 4.188459
24 16,777,216 64 1.60e--05 4.188612
28 268,435,456 128 2.00e--06 4.188651
Table 2: As $N$ varies for $f(x)=xe^{x-1}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{3/4}W$. Each row corresponds to a value of $N$ for which $n$ has increased.
@@ 4.2 Ridge functions
All of the above choices give ridge functions bounded between 00 and $1$. They span a very wide range of QMC regularity conditions. Using $g_{\mathrm{jump}}$ makes $f$ discontinuous and it has infinite variation in the sense of Hardy and Krause for $d\geqslant 2$. Using $g_{\mathrm{kink}}$ makes $f$ have infinite variation in the sense of Hardy and Krause for $d\geqslant 3$ and yet it has been seen to be a more ‘QMC friendly’ integrand than $g_{\mathrm{jump}}$ in high dimensions because it has an asymptotically bounded mean dimension mdridge , meaning that it is dominated by low dimensional interactions. It is continuous, has one weak derivative but is not differentiable everywhere. Using $g_{\mathrm{smooth}}$ makes $f$ infinitely differentiable which is generally a very QMC friendly property. Using $g_{\mathrm{finance}}$ gives $f$ a property similar to some integrands in financial option valuation, namely an unbounded derivative where $v=-2$.
@@
Figure 1: Mean confidence interval width for the HBCI as a function of the RQMC sample size $n$. There is one panel for each of the $4$ ridge functions. Within a panel the curves are for $N=2^{2r}$ for $r=4,5,6,7,8$ from top to bottom. The smallest mean width for each $N$ is marked with a larger circle. The means are taken over all $20$ replicates and all $4$ dimensions used.
@@ 4.3 Functions with known RQMC variance
Using the known variances, we can compute the half width from Bennett’s inequality  (11) for our $\sigma^{2}_{n}$ and $R=N/n$ double it and then divide the average HBCI width of 20 independent trials by this value. The results are shown in Figure 2. For our two example functions we see that as $R=N/n$ increases (so $n$ decreases for fixed $N$) the ratio of widths decreases and approaches the theoretically expected value of $1$. At the smallest $R$ (largest $n$) the HBCI were two or three times as wide as the Bennett intervals for the discontinuous function. They were somewhat wider for the smooth function.
@@
Table 3: For the given budgets $N$, we show $n_{*}$ as the minimizer of the Bennett width (13) over $n=2^{k}$ for integer $k$ and $n_{\operatorname{opt}}$ is the value of $n$ that minimized the average width of the HBCI for the given $N$. These are shown for functions $f(x)=1\{x<1/3\}$ and $f(x)=xe^{x-1}$.
@@ 5 Discussion
In favorable cases the RQMC variance is $\tilde{O}(n^{-3})$ and then using a fixed number $R$ of replicates of point sets with $n\to\infty$ we can estimate the integral $\mu$ with a standard deviation of $\tilde{O}(N^{-3/2})$ for $N=nR$ along with an unbiased estimate of the variance of that estimate. It remains very difficult to get a confidence interval of width $\tilde{O}(N^{-3/2})$ err4qmc . Using Student’s $t$ with a small $R$ gave good results in an extensive simulation by LEcEtal24a but that success is not yet theoretically understood. For an integrand subject to known bounds, we can get a non-asymptotic confidence interval by using $R$ replicates of $n$ RQMC points using a predictable plug-in empirical Berstein confidence interval. There, if $\sigma^{2}_{n}=\Theta(n^{-\theta})$ then the optimal $n$ is $\Theta(N^{1/(\theta+1)})$ and the resulting confidence intervals have width $\Theta(N^{-\theta/(\theta+1)})$. While it remains to find a way to choose the optimal $n$ empirically, we have strong theoretical guidance from equation (14) that we can interpret as an upper bound on $n$ for the oracle which is then a reasonable upper bound for HBCIs which have to estimate the RQMC variance. When $n$ must be a small power of $2$, we can rule out many suboptimal choices.
@@
It also remains to find a good way to use RQMC in confidence sequences as opposed to confidence intervals. We might set up $R$ independent infinite sequences of RQMC points and then stop them at the first $n$ where they provide a $1-\alpha$ interval narrower than some $\epsilon$. In a task like that it makes sense to only use $n=2^{k}$ for integers $k$. That is partly because powers of two are good for scrambled Sobol’ points, but also because in RQMC it is generally advisable to study sample sizes $n$ that grow geometrically not arithmetically sobo:1998 . The idea is that if $n$ is not large enough to get a good answer, then $n+1$ is unlikely to be much of an improvement. A confidence sequence based on geometrically growing $n$ would not have to be as wide as one for arithmetically growing $n$.
For scrambled Sobol’ points it is strongly advisable to take $n=2^{k}$ because those are the best sample sizes and they can even give a better convergence rate. The Halton sequence Hal60 does not have strongly superior sample sizes and so $n$ need not be a power of two for it. The two scrambles we considered for Sobol’ points can also be applied to the Halton sequence getting $\operatorname{var}(\hat{\mu})=o(1/n)$ without requiring special sample sizes $n$ haltongain . There are however no published non-trivial settings where scrambled Halton points can obtain $\theta>2$.

### v2

@@ 3 Asymptotic confidence interval comparisons
Our confidence intervals will use $R$ IID replicates that each use $n$ RQMC points in $0,1]^{d}$. That takes $N=nR$ evaluations of $f$. Here we study how to get the narrowest confidence intervals with that budget of $N$ function evaluations.
@@
We know from Waubdy-Smith and Ramdas (2024) that the widths of their predictable plug-in EBCI approaches that of some Bennett intervals described below, as $R\to\infty$ for fixed $n$. The same is true of their HBCI. Here we study the value of $n$ that optimizes the Bennett intervals.
As mentioned above, the tilde in $\tilde{O}$ hides powers of $\log(n)$. Those powers are present in error bounds for adversarially chosen integrands but do not seem to come up in real problems Schlier (2002); Owen and Pan (2022). We will consider working models of the form
@@
when using $R$ IID observations that each have variance $\sigma^{2}_{n}$. An oracle that knew $\sigma^{2}_{n}$ could use this formula to select the best $n$ and $R$ for a confidence interval subject to a constraint $nR=N$. The predictable plug-in EBCI and the HBCI from Waubdy-Smith and Ramdas (2024) do not assume a known variance for the $Y_{i}$. However those intervals attain the same asymptotic limiting width as $R\to\infty$ that an oracle would get from Bennett’s formula.
@@
If $\theta=1$, then $H(n)$ is minimized over $1,\infty)$ at $n=1$.
@@
For scrambled Sobol’ points it is generally best to take $n$ to be a power of 2 Owen (2022), which is not reflected in Theorem 1. We incorporate that restriction into some numerical examples in Section 4.
As noted above, when $\theta=1$, the best $n$ is $n=1$. In that case RQMC is the same as MC, using one uniformly distributed point in $0,1]^{d}$. This result holds for any sampler that gets the ordinary MC rate, even if it has a favorable constant. In particular, any MC variance reduction techniques that do not change the convergence rate will not lead to using $n>1$. They will improve the asymptotic confidence interval width by reducing $\sigma_{0}$.
@@ 4 Finite sample evaluations
Here we present some finite sample evaluations of the EBCI (specifically the predictable plug-in one) and the HBCI, both using RQMC, for 95% confidence. We begin with finite $N$ and for some integrands with known RQMC variance, we find the $n$ that minimizes the interval widths from Bennett’s inequality. They are then the exact best $n$ for Bennett’s inequality which the HBCIs and EBCIs of Waubdy-Smith and Ramdas (2024) match asymptotically. After that we present finite sample results on the EBCIs and HBCIs for a variety of integrands differing in smoothness and dimension. The HBCIs are usually narrower. Finally, we compare the empirically optimal $n$ for the HBCIs to the values that optimize Bennett’s inequality. The HCBI widths approach those from Bennett’s inequality for large $R$ and we find that the optimal $n$ for HBCI tend to be the same or smaller than the ones we get from Bennett’s inequality.
For our HBCI computations, we used $\theta=1/2$. The simulations presented here were run at Betting-Paper-Simulations-for-QMC. The software we used comes from Waudby-Smith (2025) which has code to reproduce the figures in Waubdy-Smith and Ramdas (2024). It calls the HBCI and predictable plug-in EBCI algorithms from Howard et al. (2021). The scrambled Sobol’ points we used were from QMCPy Choi et al. (2025).
@@ 4.1 Semi-empirical evaluations
The digital nets of Sobol’ Sobol’ (1967) under the nested uniform scramble of Owen (1995) have a stratification property. The $n$ values of $x_{ij}$ for $i=1,\dots,n$ and any component $j=1,\dots,d$ have one value uniformly distributed in each interval $(\ell-1)/n,\ell/n)$ for $\ell=1,\dots,n$ and those $n$ values are mutually independent. This allows us to work out the finite sample RQMC variances for some illustrative examples. The random linear scramble of Matoušek (1998) combined with a digital shift attains the same variance as the nested uniform scramble, but it has some dependencies among the stratified sample values. In both cases, the $d$ vectors $(x_{1j},\dots,x_{nj})\in\mathbb{R}^{n}$ are mutually independent.
We only consider values of $n$ that are powers of $2$ which is a best practice when using Sobol’ sequences Owen (2022). The smooth error rate $\tilde{O}(n^{-3/2})$ for the RMSE is attainable along powers of $2$ but not for general $n$. An elegant argument due to Sobol’ Sobol (1998) explains why. He notes that
@@
For $d=1$ and $f$ of bounded variation in the usual sense, we get $\operatorname{var}(\hat{\mu})\leqslant V_{\mathrm{HK}}(f)^{2}/n^{2}$. This holds because $D_{n}^{*}$ for the stratified sample cannot be larger than $1/n$. For illustration we choose a function with $V_{\mathrm{HK}}(f)=1$ for which we know exactly what the RQMC variance is for $n$ a power of $2$. That integrand is $f(x)=1\{x<1/3\}$. It is constant within n1n-1 of the intervals $(\ell-1)/n,\ell/n)$. All of this function’s variation happens within just one of those intervals. By inspecting the base $2$ expansion $1/3=0.010101\cdots$ we can see that the remaining value of $f(x_{\ell})$ is $1$ or 0 and it equals $1$ with a probability that is either $1/3$ if $\log_{2}(n)$ is an odd integer or is $2/3$ if $\log_{2}(n)$ is an even integer. It follows that $\operatorname{var}(\hat{\mu}_{n})=2/(9n^{2})$.
Table 1 shows how the width minimizing value of $n$ increases with $N$ in the BVHK case with $f(x)=1\{x<1/3\}$. These are widths $W=2H^{\mathrm{Ben}}(n,N/n)$ using equation (11), not half widths. Some are larger than $1$, because the Bennett width does not reflect that we would intersect the intervals with $[0,1]$. The increase in $n$ is quite slow. By $N=1024$ the best RQMC sample size is only $n=8$ and this remains the optimal sample size up to $N=4096$. The value $N^{2/3}W$ very quickly becomes constant. For non-integer $\theta$ (not shown here) the scaled widths fluctuate within a finite range instead of approaching a constant.
@@
$\log_{2}(N)$ NN $n$ $W$ $N^{2/3}W$
00 1 1 5.02e++00 5.020114
04 16 2 7.60e-01 4.826379
07 128 4 1.90e-01 4.826379
10 1,024 8 4.75e-02 4.826379
13 8,192 16 1.19e-02 4.826379
16 65,536 32 2.97e-03 4.826379
19 524,288 64 7.42e-04 4.826379
22 4,194,304 128 1.86e-04 4.826379
25 33,554,432 256 4.64e-05 4.826379
28 268,435,456 512 1.16e-05 4.826379
Table 1: As NN varies for $f(x)=1\{x<1/3\}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{2/3}W$. Each row corresponds to a value of $N$ for which $n$ has increased.
Our second example is $f(x)=xe^{x-1}$. This function is bounded between 0 and $1$. We chose it because it is smooth enough to have $\theta=3$ while also not a polynomial or an antisymmetric function which would make the problem artificially easy for some methods. Because of the stratification, we know that
@@
Table 2 shows how the optimal sample sizes $n$ grows with $N$ for our smooth example $f(x)=xe^{x-1}$. An RQMC sample size of $n=8$ becomes optimal for N=4,096N=4{,}096 and remains optimal up to N=32,768N=32{,}768. The values $N^{3/4}W$ settle down rapidly as $N$ increases. Consistent with the results of Section 3, the values $n$ grow at a slower rate with $N$ than in the BVHK case while the widths $W$ decrease at a faster rate.
$\log_{2}(N)$ NN $n$ $W$ $N^{3/4}W$
0 1 1 4.00e++00 4.003728
4 16 2 5.17e-01 4.138086
8 256 4 6.52e-02 4.175717
12 4,096 8 8.17e-03 4.185407
16 65,536 16 1.02e-03 4.187848
20 1,048,576 32 1.28e-04 4.188459
24 16,777,216 64 1.60e-05 4.188612
28 268,435,456 128 2.00e-06 4.188651
Table 2: As NN varies for $f(x)=xe^{x-1}$, this table shows the value of $n$ that minimizes the width of Bennett’s interval, along with the minimizing width $W$ and $N^{3/4}W$. Each row corresponds to a value of $N$ for which $n$ has increased.
@@ 4.2 Ridge functions
All of the above choices give ridge functions bounded between 0 and $1$. They span a very wide range of QMC regularity conditions. Using $g_{\mathrm{jump}}$ makes $f$ discontinuous and it has infinite variation in the sense of Hardy and Krause for $d\geqslant 2$. Using $g_{\mathrm{kink}}$ makes $f$ have infinite variation in the sense of Hardy and Krause for $d\geqslant 3$ and yet it has been seen to be a more ‘QMC friendly’ integrand than $g_{\mathrm{jump}}$ in high dimensions because it has an asymptotically bounded mean dimension Hoyt and Owen (2020), meaning that it is dominated by low dimensional interactions. It is continuous, has one weak derivative but is not differentiable everywhere. Using $g_{\mathrm{smooth}}$ makes $f$ infinitely differentiable which is generally a very QMC friendly property. Using $g_{\mathrm{finance}}$ gives $f$ a property similar to some integrands in financial option valuation, namely an unbounded derivative where v=2v=-2.
@@
Figure 1: Mean confidence interval width for the HBCI as a function of the RQMC sample size $n$. There is one panel for each of the 44 ridge functions. Within a panel the curves are for $N=2^{2r}$ for $r=4,5,6,7,8$ from top to bottom. The smallest mean width for each NN is marked with a larger circle. The means are taken over all 2020 replicates and all $4$ dimensions used.
@@ 4.3 Functions with known RQMC variance
Using the known variances, we can compute the half width from Bennett’s inequality (11) for our $\sigma^{2}_{n}$ and $R=N/n$, double it, and then divide the average HBCI width of 20 independent trials by this value. The results are shown in Figure 2. For our two example functions we see that as $R=N/n$ increases (so $n$ decreases for fixed $N$) the ratio of widths decreases and approaches the theoretically expected value of $1$. At the smallest $R$ (largest $n$) the HBCI were two or three times as wide as the Bennett intervals for the discontinuous function. They were somewhat wider for the smooth function.
@@
Table 3: For the given budgets NN, we show $n_{*}$ as the minimizer of the Bennett width (13) over $n=2^{k}$ for integer $k$ and $n_{\operatorname{opt}}$ is the value of $n$ that minimized the average width of the HBCI for the given NN. These are shown for functions $f(x)=1\{x<1/3\}$ and $f(x)=xe^{x-1}$.
@@ 5 Discussion
In favorable cases the RQMC variance is $\tilde{O}(n^{-3})$ and then using a fixed number $R$ of replicates of point sets with $n\to\infty$ we can estimate the integral $\mu$ with a standard deviation of $\tilde{O}(N^{-3/2})$ for $N=nR$ along with an unbiased estimate of the variance of that estimate. It remains very difficult to get a confidence interval of width $\tilde{O}(N^{-3/2})$ Owen (2026). Using Student’s $t$ with a small $R$ gave good results in an extensive simulation by L’Ecuyer et al. (2024) but that success is not yet theoretically understood. For an integrand subject to known bounds, we can get a non-asymptotic confidence interval by using $R$ replicates of $n$ RQMC points using a predictable plug-in empirical Berstein confidence interval. There, if $\sigma^{2}_{n}=\Theta(n^{-\theta})$ then the optimal $n$ is $\Theta(N^{1/(\theta+1)})$ and the resulting confidence intervals have width $\Theta(N^{-\theta/(\theta+1)})$. While it remains to find a way to choose the optimal $n$ empirically, we have strong theoretical guidance from equation (14) that we can interpret as an upper bound on $n$ for the oracle which is then a reasonable upper bound for HBCIs which have to estimate the RQMC variance. When $n$ must be a small power of $2$, we can rule out many suboptimal choices.
@@
It also remains to find a good way to use RQMC in confidence sequences as opposed to confidence intervals. We might set up $R$ independent infinite sequences of RQMC points and then stop them at the first $n$ where they provide a $1-\alpha$ interval narrower than some $\epsilon$. In a task like that it makes sense to only use $n=2^{k}$ for integers $k$. That is partly because powers of two are good for scrambled Sobol’ points, but also because in RQMC it is generally advisable to study sample sizes $n$ that grow geometrically not arithmetically Sobol (1998). The idea is that if $n$ is not large enough to get a good answer, then $n+1$ is unlikely to be much of an improvement. A confidence sequence based on geometrically growing $n$ would not have to be as wide as one for arithmetically growing $n$.
For scrambled Sobol’ points it is strongly advisable to take $n=2^{k}$ because those are the best sample sizes and they can even give a better convergence rate. The Halton sequence Halton (1960) does not have strongly superior sample sizes and so $n$ need not be a power of two for it. The two scrambles we considered for Sobol’ points can also be applied to the Halton sequence getting $\operatorname{var}(\hat{\mu})=o(1/n)$ without requiring special sample sizes $n$ Owen and Pan (2024). There are however no published non-trivial settings where scrambled Halton points can obtain $\theta>2$.